OpenStudy (otonogold):

What is the solution? 3x+3y=10 -9x-9y=-30 I am so confused! I keep getting different answers! I need help! :(

OpenStudy (anonymous):

ok first you divide by 3 to get y alone.then subtract 3x from both sides and the solution would be y=-3x+3.3

OpenStudy (anonymous):

3x+3y=10 -9x-9y=-30 lets look at the 2nd equation -9x-9y=-30 you can divide the entire thing by -3 so after that you get 3x+3y=10

OpenStudy (anonymous):

you basically have two of the same equation

OpenStudy (anonymous):

so this problem as infinately many solutions or x=all real numbers and y=all real numbers

OpenStudy (otonogold):

so it will basically be 3x+3y=10 3x+3y=10 ?

OpenStudy (thejax):

This shows no point of intersection, both the eqns coincide

OpenStudy (anonymous):

its a simultaneous equation

OpenStudy (austinl):

Multiply the top equation by three. Which should give, 9x+9y=30 -9x-9y=-30 Added together we have, 0=0 Therefore, infinitely many solutions, \(\mathbb{R}\).

OpenStudy (otonogold):

that makes a lot more sense! Thank you so much for helping me understand !